OpenStudy (anonymous):

\[20!=24329020081766xy000\] Find x,y

OpenStudy (anonymous):

x=4 and y=0

OpenStudy (anonymous):

\[10^4/20!===>y=0\] \[9/20!====>9/59+x \text{and 0<x<=9}===>x=4\]

OpenStudy (turingtest):

I think you missed a zero @mukushla which would make @Neemo 's answer correct

OpenStudy (turingtest):

http://www.wolframalpha.com/input/?i=20!&t=crmtb01

OpenStudy (anonymous):

4 0's at the end and \[\frac{20!}{10^4} \equiv\ 4 (mod \ 10)\]

OpenStudy (turingtest):

oh wait, I may have miscounted the digits... sorry

OpenStudy (anonymous):

np :)